OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

4x - 3y = -12 4x+12=3y i.e. y= 4/3x+4, comparing it with y=mx+c we find: slope m= 4/3 Parallel lines have same slopes Therefore, slope of required line m= 4/3 since line passes through point (-3, 4) Equation of line in slope point form is given as y-y1=m(x-x1) y-4=4/3(x-(-3)) 3(y-4)=4(x+3)) 3y-12=4x+12 3y=4x+24 4x-3y+24=0 is the required equation of the line @greenlegodude57